A 95kg clock is pushed along the floor at constant speed by a force of 560N . Calculate the coefficient of friction.
1
Expert's answer
2020-01-27T09:31:12-0500

Applying the Newton’s Second Law of Motion we get (since the clock is pushed along the floor at constant speed the acceleration is zero):


"\sum {F_{x}} = ma_x = 0,""F_{push} - F_{fr} = 0,""F_{push} = F_{fr} = \mu N = \mu mg."

From this formula we can find the coefficient of friction:


"\mu = \dfrac{F_{push}}{mg} = \dfrac{560N}{95kg \cdot 9.8 \dfrac{m}{s^2}} = 0.60"

Answer:

"\mu = 0.60"